A firm projects an ROE of 10% and it will maintain a plowback (reinvestment) ratio of 0.2. The firm is expecting earnings of R4 per a share and the investors expect a return 10% on the stock. Calculate the expected price and P/E ratio of the firm.
Answer #1
Growth rate=ROE*retention ratio
growth rate=10*0.2
growth rate = 2

DPS = EPS*(1-plowback) = 4*(1-0.2) = 3.2

Price = Dividend in 1 year/(cost of equity - growth rate)
Price = 3.2/ (0.1 - 0.02)
Price = 40

P/E = price/EPS = 40/4 = 10